# If quasars form from primordial black holes

### v1

@@ Abstract
We explore the consequences of a novel but increasingly well-supported hypothesis that supermassive black holes may have formed from primordial black holes formed prior to, and rapidly growing in, the radiation-dominated universe. We show that this hypothesis can predict the luminosity of quasars and their luminosity distribution. With reasonable values of the parameters introduced, these predictions are borne out by observations. The model predicts density evolution in accordance with observations. If the same galaxy interaction rate creates quasars and radio galaxies, whose primordial black hole nuclei seem somewhat less massive, their relative number densities reflect relative lifetimes in these states.
@@ 1 Introduction
Two breakthroughs, their discovery (Schmidt 1963) and identification of their black hole engines (Lynden-Bell & Rees 1971), marked the arrival of a new era of higher redshift astronomy. Their physical understanding has been one of astronomy’s greatest challenges. This is because the QSOs’ efficiency in energy production causes them to employ every conceivable channel for energy transfer right across the electromagnetic spectrum, plus radiation driven mass loss. A review for by Antonucci (2013) claimed ”Quasars still defy explanation”, and, while much knowledge has been added, that statement remains true today.
A different way of looking at the physics is to consider their fuelling. This has worked well with stellar evolution, but, as in stellar evolution, it tells us about the numbers of objects at different stages, not, for example, details of their spectral energy distribution, which is a product of much more involved physics, see e.g. Porras-Valverde et al. ( 2025). Here we consider their possible formation as primordial black holes (PBH 2) (Bicknell & Henricksen 1979; Volonteri, Habouzit & Colpi 2021; Davies, Miller & Bellovary, 2011, Lupi et al. 2014; Ziparo, Gallerani & Ferrara 2025; Cai et al. 2024; Dayal & Maiolino 2025; Sobrinho & Augusto 2024) during the radiation dominated era (RE) of the universe (Mould 2025). Parameterizing variables that we don’t know, we find we can successfully conform the model to their average bolometric luminosity (2) and the luminosity function in 3. In 4 we draw attention to the imprint of inflation on the origin of the seeds of SMBH, and conclude that the model can be further tested with galaxy formation simulations and large volume surveys for QSOs with JWST, and also the forthcoming space telescope and other observations.
@@ 2.1 Initial conditions
According to the Friedmann equation the density during the RE is given by 1/ = 32/3 or r = ct/8, where t is the age of the universe and r is the Schwarzschild radius of the primordial black hole that would form. There is an extensive literature on the seeds of PBHs originating during inflation. Reviews are given by Ögsoy & Tasinato (2023) and Harada (2024). This RE radius equation (which neglects the density excess that is required to actually form a PBH111r = ct/8 is not exact, see, for example, Germani & Sheth (2023). The mass excess inside a given volume is called a compaction function C(r). The peak value of the compaction function is used as a criteria for PBH formation. Simulations show the existence of a lower bound C 2/5; there is also an upper bound C 2/3. The precise value of the lower bound is not tightly determined, but recently it has been shown by numerical studies that slightly lower values are possible, and that the threshold lies in the range 1/5 C 1/3. Atal et al. (2020) give a shape-independent threshold value for gravitational collapse, C 1/5. For now we simply take 1+C as a constant of order unity, as it is degenerate with the parameter f in equation (1). We return to note the implication of this correction in 2.4 ) allows PBH as large as 10 million M⊙ to form at t = 400 seconds at the end of primordial nucleosynthesis. We assume that a fraction f of the dark matter in the Universe is in the form of PBH of mass 10-8 M⊙ (denoted m8) with an initial mass function (IMF) that places equal mass in each decade222A local IMF, not one that stretches very many orders of magnitude. PBH formation may be associated with well separated phase transitions (Kuhnel, Rampf & Sandstad 2016). of mass (Mould 2025), n m-1. Given the density of dark matter from the Planck collaboration (2020), we can write for the number density of m8
@@
Figure 1 shows two different IMFs for PBHs, together with an approximate constraint on their numbers from microlensing. A constraint on more massive PBH from LIGO is similar for masses between 10 and 300 M⊙ (Ali-Haïmoud, Kovertz & Kamionkowski 2017).
@@ 2.3 Fuel burning
This collision rate is for SMBH and 10-8 M⊙ PBH, but the identification with dwarf galaxies is made for dwarf galaxy sized clumps of dark matter, and the gaps in time between them correspond to QSO inactive periods. This standard collision rate formula is referenced in the Appendix. Equation (2) is therefore appropriate for SMBH activated by dwarf accretion. Orbital decay of the dwarf may be slow (Conselice et al. 2019). we consider this part of the inactive period. Supposing 100% conversion of mass (all of the dwarf is consumable333Maximum values of between 0.1 and 0.4 are considered possible (Blandford & Begelman 1999; Kovacs et al. 2011), baryons and dark matter), the luminosity density of quasars is
@@
and a range of 0.5 dex in M7 would give a range of 1 dex in luminosity. Shen et al. (2020) find this to be = –4.7 cMpc-3/dex for L = L46 at z = 1.
@@ 2.4 Parameter constraints
for b 1.
A value of b and v200 can be obtained from simulations by comparing the actual interaction rate of galaxies with the reaction rate with geometric cross sections. The halo mass function would also be involved. In the nearby SMBH sample of Winkel et al. (2025) 2, but the 4 objects mentioned in the database NED444ned.ipac.caltech.edu, 3C273, PG0026+129, Mrk1501, PG1211+143 as QSOs, as opposed to Seyfert galaxies, have a mean M7 of 23 and a dispersion of 0.5 dex. At present microlensing experiments provide the best way of estimating f. Our setting the PBH compaction function to one instead of the current best estimate, one fifth, has increased the predicted QSO luminosity density by a factor of 1.46. Correction is degenerate with the parameters f, and b.
@@ 2.5 Intermediate mass black holes
IMBHs remain a puzzle in astrophysics because M 104 M⊙ black holes cannot be stellar remnants (Davis et al. 2024), unless perhaps in galaxies earlier than those we can observe, in Population III. Formation as PBHs is, however, possible in principle. If IMBH are confirmed in galactic or globular cluster nuclei, for example, it would be a further demonstration of a role for PBHs, in addition to the involvement in QSO formation we are discussing here.
@@ 3 The quasar luminosity function
The previous section gave us the expected number density of quasars accreting in this way. We combine this with the PBH IMF for the SMBH PBHs from Mould (2025), log n + log M = constant, and with a little algebra555 From equation (3) n(L) = for L and dlogn = -2 dlogM and, integrating over the IMF, log n = 2 1/m dlogm., we obtain Figure 2. For this PBH IMF, the form of the curve from the previous section is obtained by integrating over it to get
where x = log L, x′ = x - 46 and L42 is the luminosity of the minimum mass PBH progenitor, accounting for (42 - x) in the exponent. This luminosity function is a Schechter function (power law with exponential cut off, Schechter 1976). It would be useful to find out if this is a better or worse fit to actual data than a double power law666A double power law can be formed from eqn (6) by bisecting it at L46.. There is one less degree of freedom in the single power law and (x0 = 42, y0 = -4.7) are fully independent, the y coordinate being the y-axis of Figure 2. The value of the minimum luminosity QSO, four orders of magnitude below L46, points directly to the post RE mass of its SMBH progenitor which would be a 106 M⊙ PBH. ’Minimum luminosity’ should be regarded as a fitting parameter, rather than a physical constraint. Furthermore, Mould (2025) shows that the cut-on mass of the IMF is a sharp marker of the cosmic temperature. The excellent fit777Shen et al. (2020) fits to the data at z = 1. in black, model in green of the log M + log n = constant IMF is an indicator that density inhomogeneities with n M-1 created during inflation were responsible for the PBH IMF.
In the paradigm of 2 it is interesting to ask why there is a QSO luminosity function (LF) rather than a delta function at L46. And is the mean QSO luminosity a standard candle ? Two prime candidates for sources of a distribution function are the m7 and M7 parameters in the model. The PBH IMF creates spreads in both parameters. If only these parameters have a range, density evolution alone would be expected. A diverse formation history of dwarf galaxies across the relevant redshift range would lead to luminosity evolution as well. For = 1 a fuelling rate of 10 per 108 years (a free fall time) produces 6 1046 ergs/s. The astrophysics that modulates this luminosity on shorter timescales is certainly complex, and a subject in itself. But it is the economics of supply that was considered in 2. If the overall efficiency of consumption of galaxies by quasars were, say, 0.01%, this could be accommodated in equation (5) by an order of magnitude change in M7, the principle would still hold, and the LF would remain as shown.
@@ 3.1 Luminosity and density evolution
Equation (4) gives the dependence of quasar luminosity density on the redshift of the observer, = –3. Figure 3 shows density evolution versus the scale factor. A slope of -3 is expected from the model, and this is close to the data. Luminosity evolution would require a variation of one or more of the parameters with redshift to remain within the framework of the model. Both m8 (the fuel) and M7 (the consumer) are candidates. Since the mass of the SMBH or its mean value was laid down at inflation, it is effectively ruled out as a variable across the observable redshift range. The supply of dwarf galaxies as fuel, however, is plausible as the responsible factor. Observations can test this paradigm, both of galaxy number density versus mass and redshift and of QSOs at JWST redshifts (Labbé et al. 2025) to extend Figure 2. The large error bar on the z 10 point in Figure 2 allows the pure density evolution of the model to fit the data. But a future square degree sample of 100 or more QSOs, confirmed with spectra, would shrink the errors bars to 0.05 dex. Matthee et al. (2024) measure number densities of 10-5 cMpc-3. Kocevski et al. (2024) find that ”little red dots” are 100–1000 more numerous than bright quasars at z 5–7, but their number density is only 4–10 times higher than X-ray and UV selected AGN.
If the present consensus is confirmed that pure density models do not fit the abundance of AGN at redshift 10, these models will suggest different spectral energy distributions and bolometric corrections or faster buildup of SMBH. However, fuelling models should also be considered with, perhaps , even an extreme duty cycle with QSOs ”always on”. And, although m8 PBH do not lose mass during galaxy formation, m17–m19 PBH do, and would have been able to supply more fuel at z = 10.
Shen et al. (2020) present a careful analysis of the bolometric LF variation with redshift, using all the data rather than the coarse bins illustrated here. They find both luminosity and density evolution. The key question for large volume JWST ”little red dot” surveys is, does the naive version of our model, pure density evolution, still fit, or, are the parameters evolving in some way that sheds light on QSO power?
@@ 3.2 A QSO standard candle
There is an extensive cosmology literature in which quasars are used as standard candles (e.g. Signorini et al. 2024). Figure 2 is supportive of this assumption, although the pure density evolution slope is quite uncertain without higher redshift data. Minimum, mean888The mean luminosity for equation (6) is 4 L46. and maximum standard luminosities and their dispersion are afforded by the model outlined in 2 as a function of redshift. With higher redshift data arriving from JWST the wealth of data can be expected to amply constrain the simple model presented here. Traditionally stars and supernovae are seen as standard candles, and galaxies with their complex history less so. Quasars have their roots in PBH which follow a simple law, although accretion may complicate it. This is a physical basis for QSOs as standard candles.

### v2

# Do quasars form from primordial black holes ?

@@ Abstract
We explore the consequences of a novel but increasingly well-supported hypothesis that supermassive black holes may have formed from primordial black holes (PBH) formed prior to, and rapidly growing in, the radiation-dominated universe. We show that this hypothesis and fuelling by accretion of dwarf galaxies, made up of predominantly subsolar mass PBH and accompanying gas, can predict the luminosity of quasars and their distribution in luminosity. With reasonable values of the parameters introduced, these predictions are borne out by observations. The model predicts density evolution in accordance with observations. If the same galaxy interaction rate creates quasars and radio galaxies, whose primordial black hole nuclei seem somewhat less massive, their relative number densities reflect relative lifetimes in these states.
Primordial black holes(1292) - Cosmology(343) - Quasars(1319) - Supermassive black holes(1663) - Radio galaxies(1343)
@@ 1 Introduction
Two breakthroughs, their discovery (Schmidt 1963) and identification of their black hole engines (Lynden-Bell & Rees 1971), marked the arrival of a new era of higher redshift astronomy. Their physical understanding has been one of astronomy’s greatest challenges. This is because the QSOs’ efficiency in energy production causes them to employ every conceivable channel for energy transfer right across the electromagnetic spectrum, plus radiation driven mass loss. A review for by Antonucci (2013) claimed "Quasars still defy explanation", and, while much knowledge has been added, that statement remains true today.
A different way of looking at the physics is to consider their fuelling. This has worked well with stellar evolution, but, as in stellar evolution, it tells us about the numbers of objects at different stages, not, for example, details of their spectral energy distribution, which is a product of much more involved physics, see e.g. Porras-Valverde et al. ( 2025). Here we consider their possible formation as primordial black holes (PBH 2) (Bicknell & Henricksen 1979; Volonteri, Habouzit & Colpi 2021; Davies, Miller & Bellovary, 2011, Lupi et al. 2014; Belotsky et al. 2019; Ziparo, Gallerani & Ferrara 2025; Cai et al. 2024; Dayal & Maiolino 2025; Sobrinho & Augusto 2024; Li, Huang & Piao 2025, Zhang et al. 2025; Hooper, Ireland, & Krnjaic 2025; Liu & Bromm 2025; Prole & Regan 2025 ) during the radiation dominated era (RE) of the universe (Mould 2025). Parameterizing variables that we don’t know, we find we can successfully conform the model to their average bolometric luminosity (2) and the luminosity function in 3. In 4 we draw attention to the imprint of inflation on the origin of the seeds of SMBH, and conclude that the model can be further tested with galaxy formation simulations and large volume surveys for QSOs with JWST, and also the forthcoming space telescope and other observations.
@@ 2.1 Initial conditions
According to the Friedmann equation the density during the RE is given by 1/ = 32/3 or r = ct/8, where t is the age of the universe and r is the Schwarzschild radius of the primordial black hole that would form. There is an extensive literature on the seeds of PBHs originating during inflation. Reviews are given by Ögsoy & Tasinato (2023) and Harada (2024). This RE radius equation (which neglects the density excess that is required to actually form a PBH) allows PBH as large as 10 million M⊙ to form at t = 400 seconds at the end of primordial nucleosynthesis. In addition to supposing SMBH are PBH, we assume that a fraction f of the dark matter in the Universe is in the form of PBH of mass 10-8 M⊙ (denoted m8) with an initial mass function (IMF) that places equal mass in each decade111A local IMF, not one that stretches very many orders of magnitude. PBH formation may be associated with well separated phase transitions (Kuhnel, Rampf & Sandstad 2016). of mass (Mould 2025), n m-1. Given the density of dark matter from the Planck collaboration (2020), we can write for the number density of m8
@@ 2.2 CMB distortions
2.3 PBH IMF
Kühnel, Rampf & Sandstad (2016) studied the formation of PBH in several models of the early Universe. PBH production has generally been approximated to yield black holes of horizon size at the time of formation. However, it was found and subsequently confirmed in several thorough numerical works that the PBHs are formed through critical collapse. The critical collapse leads to a spectrum for the formation of PBHs at any time over a wide range of masses, including 10-11 to 105 M⊙. Here we adopt a minimalist assumption that equal masses exist in decades of mass (n m-1) around particular peaks, exemplified by Kühnel, Rampf & Sandstadt’s figure 4, and, for present purposes, we concentrate (1) on subsolar masses making up the dominant DM in dwarf galaxies, and (2) SMBH. The mass cut-offs in the subsolar regime can be adopted from microlensing observations: the "asteroid window" ranges from 10-11 to 10-7 M⊙ (Niikura et al. 2019; Sugiyama, Takada, Yasuda & Tominaga 2026). We return to this subject in the next section, and to the supermassive range in 2.7.
@@ 2.4 Fuel burning
This collision rate is for SMBH and 10-8 M⊙ PBH, but the identification with dwarf galaxies is made for dwarf galaxy sized clumps of dark matter, and the gaps in time between them correspond to QSO inactive periods. This is a standard collision rate formula (Clayton 1968). Equation (2) is therefore appropriate for SMBH activated by dwarf accretion. Orbital decay of the dwarf may be slow (Conselice et al. 2020). we consider this part of the inactive period. Supposing 100% conversion of mass (all of the dwarf is consumable222Maximum values of between 0.1 and 0.4 are considered possible (Blandford & Begelman 1999; Kovacs et al. 2011), baryons and dark matter), the luminosity density of quasars is
@@
and a range of 0.5 dex in M7 would give a range of 1 dex in luminosity. Shen et al. (2020) find this to be = –4.7 cMpc-3/dex for L = L46 at z = 1, compared with –3.8 for f = K = b = 1. Choosing to normalize at z = 2 would change this number to –4.8, compared to –5.0 from Shen et al. .
@@ 2.5 Parameter constraints
for b 1. Gravitational focussing can inform the value of b, and suggests b = 1 + (v. For a SMBH this is GM/R/, where R is the radius of the accretion disk,. 1 + M/10 1, where is the surrounding velocity dispersion in 30 km s-1unitsi and M is in solar units.. For two compact objects, bound capture requires some form of dissipation, and this reduces the term, but not in a ready physically quantifiable way. If dissipation took 50% of the energy per unit mass, , for an M7 SMBH, b = 1 + 200 (30/200)2 = 5.5 for = 200 km s-1.
A value of b and v200 can be obtained from simulations by comparing the actual interaction rate of galaxies with the reaction rate with geometric cross sections. The halo mass function would also be involved. In the nearby SMBH sample of Winkel et al. (2025) 2, but the 4 objects mentioned in the database NED333ned.ipac.caltech.edu, 3C273, PG0026+129, Mrk1501, PG1211+143 as QSOs, as opposed to Seyfert galaxies, have a mean M7 of 23 and a dispersion of 0.5 dex. At present microlensing experiments provide the best way of estimating f. Our setting the PBH compaction function to one instead of the current best estimate, one fifth, has increased the predicted QSO luminosity density by a factor of 1.46. Correction is degenerate with the parameters f, and b.
2.6 Physical Context of the accretion and fuelling
QSOs show a characteristic big blue bump feature in the UV (Mandal, Woo & Wang 2025), with the total luminosity at 1m typically 20%-40% of that at 1m. In this model the emission in the spectral range 10–300 nm, the "big blue bump, is dominated by 10,000-100,000 K thermal emission from an accretion disk. The emission between 2 microns and 1 mm, the "infrared bump," is made up of reradiation from dust in a distorted disk extending from 0.1 pc to more than 1 kpc (Sanders, Phinney, Neugebauer, Soifer & Matthews 1989). The UV radiation emanates from a concentrated region surrounding the SMBH, which forms a geometrically thin, optically thick accretion disk, which emits thermal radiation in the X-ray to optical wavelength range with a peak in the UV (Shakura & Sunyaev 1973). This is a vital component of our model, and the crux of the accretion scenario, but we deal here only with the bolometric output.
@@ 2.7 Intermediate mass black holes
In the paradigm in which there are only stellar mass black holes and SMBH, intermediate mass black holes (IMBH) are a conundrum. They arise naturally from the present PBH hypothesis, however. A second issue with IMBH is observational (Greene, Strader, & Ho 2020). The stellar mass, SMBH mass relation becomes sparse at 105 M⊙. This, and the uncertainty around globular cluster nuclei, is most likely an observational selection effect, whose cure is improved adaptive optics, to get close to the nucleus. The mass determined for a nuclear black hole is v2r/G, where v is the velocity width of an emission line and r is the distance from the nucleus of the ring of emitting gas. The smaller r, the further into IMBH territory the data require the mass derived to be. IMBHs remain a puzzle in astrophysics because M 104 M⊙ black holes cannot be stellar remnants (Davis et al. 2024), unless perhaps in galaxies earlier than those we can observe, in Population III. Formation as PBHs is, however, possible in principle. If IMBH are confirmed in galactic or globular cluster nuclei, for example, it would be a further demonstration of a role for PBHs, in addition to the involvement in QSO formation we are discussing here.
@@ 3 The quasar luminosity function
The previous section gave us the expected number density of quasars accreting in this way. We combine this with the PBH IMF for the SMBH PBHs from Mould (2025), log n + log M = constant, and with a little algebra444From equation (3) n(L) = for L and dlogn = -2 dlogM and, integrating over the IMF, log n = 2 1/m dlogm., we obtain Figure 1. For this PBH IMF, the form of the curve from the previous section is obtained by integrating over it to get
where x = log L, x′ = x - 46 and L42 is the luminosity of the minimum mass PBH progenitor, accounting for (42 - x) in the exponent. This luminosity function is a Schechter function (power law with exponential cut off, Schechter 1976). It would be useful to find out if this is a better or worse fit to actual data than a double power law555A double power law can be formed from eqn (6) by bisecting it at L46.. There is one less degree of freedom in the single power law and (x0 = 42, y0 = -4.7) are fully independent, the y coordinate being the y-axis of Figure 1. The value of the minimum luminosity QSO, four orders of magnitude below L46, points directly to the post RE mass of its SMBH progenitor which would be a 106 M⊙ PBH. ’Minimum luminosity’ should be regarded as a fitting parameter, rather than a physical constraint. Furthermore, Mould (2025) shows that the cut-on mass of the IMF is a sharp marker of the cosmic temperature. The excellent fit666Shen et al. (2020) fits to the data at z = 1. in black, model in green of the log M + log n = constant IMF is an indicator that density inhomogeneities with n M-1 created during inflation were responsible for the PBH IMF.
3.1 Parameter sensitivity
The normalization of Figure 1 is equivalent to setting f = 0.125 in equation (4). This is a plausible value for the fraction of the DM that is PBH, given the microlensing constraints. But f is degenerate with K and b in equation (4), and they have the same effect on Figure 1: moving the black curve up and down. If b = 5, then f = 0.025 would restore the fit.
In the paradigm of 2 it is interesting to ask why there is a QSO luminosity function (LF) rather than a delta function at L46. And is the mean QSO luminosity a standard candle ? Two prime candidates for sources of a distribution function are the m8 and M7 parameters in the model. The PBH IMF creates spreads in both parameters. If only these parameters have a range, density evolution alone would be expected. A diverse formation history of dwarf galaxies across the relevant redshift range would lead to luminosity evolution as well. For = 1 a fuelling rate of 10 per 108 years (a free fall time) produces 6 1046 ergs/s. The astrophysics that modulates this luminosity on shorter timescales is certainly complex, and a subject in itself. But it is the economics of supply that was considered in 2. If the overall efficiency of consumption of galaxies by quasars were, say, 0.01%, this could be accommodated in equation (5) by an order of magnitude change in M7, the principle would still hold, and the LF would remain as shown.
@@ 3.2 Luminosity and density evolution
Equation (4) gives the dependence of quasar luminosity density on the redshift of the observer, = –3. Figure 2 shows density evolution versus the scale factor. A slope of -3 is expected from the model, and this is close to the data. Luminosity evolution would require a variation of one or more of the parameters with redshift to remain within the framework of the model. Both m8 (the fuel) and M7 (the consumer) are candidates. Since the mass of the SMBH or its mean value was laid down at inflation, it is effectively ruled out as a variable across the observable redshift range. The supply of dwarf galaxies as fuel, however, is plausible as the responsible factor. Observations can test this paradigm, both of galaxy number density versus mass and redshift and of QSOs at JWST redshifts (Labbé et al. 2025) to extend Figure 2. The large error bar on the z 10 point in Figure 2 allows the pure density evolution of the model to fit the data. But a future square degree sample of 100 or more QSOs, confirmed with spectra, would shrink the errors bars to 0.05 dex. Matthee et al. (2024) measure number densities of 10-5 cMpc-3. Kocevski et al. (2024) find that "little red dots" are 100–1000 more numerous than bright quasars at z 5–7, but their number density is only 4–10 times higher than X-ray and UV selected AGN.
If the present consensus is confirmed that pure density models do not fit the abundance of AGN at redshift 10, these models will suggest different spectral energy distributions and bolometric corrections or faster buildup of SMBH. However, fuelling models should also be considered with, perhaps , even an extreme duty cycle with QSOs "always on". And, although m8 PBH do not lose mass during galaxy formation, m17–m19 PBH do, and would have been able to supply more fuel at z = 10.
Shen et al. (2020) present a careful analysis of the bolometric LF variation with redshift, using all the data rather than the coarse bins illustrated here. They find both luminosity and density evolution. The key question for large volume JWST "little red dot" surveys is, does the naive version of our model, pure density evolution, still fit, or, are the parameters evolving in some way that sheds light on QSO power?
@@ 3.3 A QSO standard candle
There is an extensive cosmology literature in which quasars are used as standard candles (e.g. Signorini et al. 2024). Figure 2 is supportive of this assumption, although the pure density evolution slope is quite uncertain without higher redshift data. Minimum, mean777The mean luminosity for equation (6) is 4 L46. and maximum standard luminosities and their dispersion are afforded by the model outlined in 2 as a function of redshift. With higher redshift data arriving from JWST the wealth of data can be expected to amply constrain the simple model presented here. Traditionally stars and supernovae are seen as standard candles, and galaxies with their complex history less so. Quasars have their roots in PBH which follow a simple law, although accretion may complicate it. This is a physical basis for QSOs as standard candles.
Risaliti et al. (2023) find a Hubble-Lemaitre diagram with redshifts out to 5 that differs from the CDM predictions by 4. The standard candle is from an ultraviolet-xray luminosity relation that is non-linear (logarithmic slope 0.63). There is no theoretical basis for this relation, and our model does not provide one, as the spectral energy distribution is not modelled here. While equation (4) from fuelling considerations does imply a bolometric luminosity standard candle, and the parameters, b and f are unlikely to evolve with redshift, the parameter K involves complex astrophysics, and SMBH mass is measureable.